The brain volumes ​(cm3​) of 20 brains have a mean of 1161.2cm3 and a standard deviation of122.7 . Use the given standard deviation and the range rule of thumb to identify the limits separating values that are significantly low or significantly high. For such​ data, would a brain volume of 1446.6 be significantly​ high?


1
Expert's answer
2020-09-10T19:30:30-0400

Mean = 1161.2

Standard deviation = 122.7

The range rule of thumb indicates that the usual values are within 2 standard deviation from the mean.

Hence lower limit = Mean - 2 x Standard deviation

=1161.2 - 2x122.7

=1161.2 - 245.4

=915.8

Upper Limit = Mean + 2 x Standard deviation

=1161.2 + 2x122.7

=1161.2 + 245.4

=1406.6

The limit of values is from 915.8 to 1406.6


A value below 915.8 is considered low and a value above 1406.6 is considered high


1446.6 is more than the upper limit. Thus it is high for a brain.